OpenStudy (anonymous):

HELP PLEEEAAASSSSEEEE !!!!!!!!!!!!!!!! Write the equation of the line parallel to x – y = 4 and passing through the point (1,6).

OpenStudy (anonymous):

The parallel function to y = x - 4, passing through (1,6) is y = x + 5